### v1

# Permanental Analog of the Rank-Nullity Theorem for Symmetric Matrices

Abstract
The rank-nullity theorem is a core result in the study of matrices. The rank of an $n\times n$ matrix is equal to the size of its largest square submatrix with a nonzero determinant; it can be computed in $O(n^{2.37})$ time. For symmetric matrices, the rank equals to the number of nonzero eigenvalues (that is, the nonzero roots of the characteristic polynomial), which implies that the nullity equals the multiplicity of zero eigenvalue. Similar to the rank, the permanental rank of a matrix is the size of the largest square submatrix with nonzero permanent (whose computation is #P-complete), while the permanental nullity is the multiplicity of zero as a root of the permanental polynomial. In this paper, we prove a permanental analog of the rank-nullity theorem for any non-negative symmetric matrix, positive semi-definite matrix and (-1, 0, 1)-balanced symmetric matrices.
1 Introduction
The determinant of an $n\times n$ matrix $A=(a_{ij})$ is defined as
$\det(A)=\sum_{\sigma\in S_{n}}\operatorname{sgn}(\sigma)\prod_{i=1}^{n}a_{i,\sigma(i)},$
where $S_{n}$ is the set of all permutations of $\{1,2,\ldots,n\}$ and $\operatorname{sgn}$ is the sign of the permutation $\sigma$. The permanent is defined in a similar way, but without the $\operatorname{sgn}$ factor:
$\operatorname{per}(A)=\sum_{\sigma\in S_{n}}\prod_{i=1}^{n}a_{i,\sigma(i)}.$
Although the definition differs only by a sign factor, the computational complexity of the determinant and permanent varies significantly. The determinant can be computed in time $\mathcal{O}(n^{2.37})$ [12], while finding the permanent is known to be #$\#P$-complete [27]. The permanent lacks a well established algebraic or geometric interpretation, and it is neither multiplicative nor invariant under linear combinations of rows or columns. As a result, the permanent received limited attention in the early literature, with nearly all known results at the time compiled in the book [19]. In 1979, Valiant [27] proved that computing the permanent is #$\#P$-complete, thus making the permanent a central object of study in computational complexity theory. For more work on the complexity and structural properties of the permanent and determinant, see [22].
The Pólya Permanent Problem [21] investigates conditions under which the permanent of a given matrix can be transformed into the determinant of a modified matrix. This problem has been shown to be equivalent to twenty-three other combinatorial and graph-theoretic problems [17], including the enumeration of perfect matchings in bipartite graphs. This motivates the search for algebraic frameworks in which the permanent satisfies identities structurally similar to those of the determinant.
Unlike the determinant, the permanent generally lacks multiplicativity, that is, $\operatorname{per}(AB)\neq\operatorname{per}(A)\operatorname{per}(B)$. Marcus and Minc [16] conjectured that multiplicativity holds only when both matrices are products of permutation and diagonal matrices. Beasley [3] later proved this, showing it is the maximal class for which the permanent is multiplicative. Beyond multiplicativity, several determinant inequalities have been adapted for the permanent. Lieb [14] established a reversed version of the classical Fischer inequality for permanents. Specifically, for a block matrix $A=\begin{pmatrix}B&C\\ C^{*}&D\end{pmatrix}\succeq 0,$ where $A$ is positive semidefinite, the inequality $\operatorname{per}(A)\geq\operatorname{per}(B)\operatorname{per}(D)$ holds. This stands in contrast to the determinant case, where the inequality is reversed, that is, $\det(A)\leq\det(B)\det(D)$ under the same conditions. Carlen, Lieb, and Loss also proved a Hadamard type bound, relating the permanent to the product of the row norms [5].
Heuvers, Cummings, and Bhaskara Rao [11] established that the permanent satisfies a Binet–Cauchy-type identity analogous to that of the determinant. For an $m\times n$ matrix $A$ be an $m\times n$ matrix and $B$ an $n\times m$ matrix with $m\leq n$. Then,
$\operatorname{per}(AB)=\sum_{k}\frac{1}{\mu(k)}\operatorname{per}[a_{k_{1}},\dots,a_{k_{m}}]\operatorname{per}\langle b_{(k_{1})},\dots,b_{(k_{m})}\rangle,$
where the sum is over all weakly increasing $m$-tuples of integers $1\leq k_{1}\leq\cdots\leq k_{m}\leq n$. $a_{k_{i}}$ denotes the $k_{i}$-th column of $A$, and $b_{(k_{i})}$ denotes the $k_{i}$-th row of $B$. The matrix $[a_{k_{1}},\dots,a_{k_{m}}]$ is the $m\times m$ matrix formed by selecting the specified columns of $A$, while $\langle b_{(k_{1})},\dots,b_{(k_{m})}\rangle$ is the $m\times m$ matrix formed by selecting the specified rows of $B$. The quantity $\mu(k)=\prod_{i=1}^{n}v_{i}(k)!$ where $v_{i}(k)$ denotes the number of times the index $i\in[n]$ appears in the tuple $k$.
Analogous to the characteristic polynomial, the permanental polynomial of an $n\times n$ matrix $A$ is defined as
$\pi(A,x)=\operatorname{per}(xI-A),$
where $I$ is the identity matrix of order $n$. Given a graph $G$ with adjacency matrix $A(G)$, its permanental polynomial is $\pi(G,x)=\mathrm{per}(xI-A(G))$. The multiset of all roots of $\pi(G,x)$, including multiplicities, is called the per-spectrum of $G$. The study of permanental polynomials in graph theory began with Turner [26], and was later extended by Merris et al. [18] and Kasum et al. [13], who explored their mathematical and chemical relevance. For results on relationship between rank, determinant, and permanent in structured graphs, see [23]. For more results on permanental polynomials and spectra, see [6, 9, 15, 32, 4, 24, 25, 2, 7].
The permanental nullity, $\eta_{\text{per}}(G)$ defined as the multiplicity of zero as a root in $\pi(G,x)$ was first studied by Wu and Zhang [28]. They established its connection to the matching number via the Gallai–Edmonds structure theorem. They provided exact characterizations for graphs with extremal per-nullities $n-2,n-3,n-4,n-5$, and derived a sharp formula for general graphs involving factor-critical components. Their work also characterized graphs with zero permanental nullity and analyzed its behavior on unicyclic graphs, line graphs, and factor-critical graphs.
The concept of permanental rank, denoted by $\rho_{\text{per}}(A)$ for a matrix $A$, was introduced by Yang Yu [29] as a combinatorial analogue of the classical matrix rank. Given a matrix $A\in\mathbb{F}^{n\times m}$, $\rho_{\operatorname{per}}(A)$ is defined as the maximum integer $k$ such that $A$ contains a $k\times k$ submatrix whose permanent is nonzero.
Yu established the following fundamental inequality relating permanental rank and classical rank: $\operatorname{rank}(A)\leq 2\rho_{\operatorname{per}}(A),$ which is shown to be tight. Furthermore, over a field of positive characteristic $p>0$, they proved that if a matrix has a column repeated $p$ times, then its permanent vanishes. This leads to an upper bound of the form: $\rho_{\operatorname{per}}(A)\leq(p-1)\operatorname{rank}(A).$
Motivated by the Alon–Jaeger–Tarsi Conjecture [1], they proposed a stronger conjecture: for any nonsingular $n\times n$ matrix $A$, the permanental rank of the matrix $(A\mid A)\in\mathbb{F}^{n\times 2n}$, formed by horizontally concatenating $A$ with itself, satisfies $\rho_{\operatorname{per}}(A\mid A)=n.$ This has been verified for small dimensions and fields of characteristic $p=3$. Additionally, for any integer $l$ and nonsingular matrices $A_{1},A_{2},\dots,A_{\ell}\in\mathbb{F}^{n\times n}$, they proved the following bound:
$\rho_{\operatorname{per}}(A_{1}A_{2}\cdots A_{\ell})\geq\left(1-\frac{2}{\ell}\right)n,$
though the tightness of this inequality is not yet known. In a probabilistic setting, they further showed that for a uniformly random nonsingular matrix $A$ over a finite field, the probability that $\rho_{\operatorname{per}}(A\mid A)=n$ tends to $1$ as $n\to\infty$, thus providing asymptotic support for the conjecture.
Fanaï [8] explored the connection of $\rho_{\operatorname{per}}(A)$ with the existence of partial transversals in additive Latin squares, particularly those arising from group-theoretic structures like Cayley tables.
In this paper, we use the terms cycle cover and partial cycle cover in the context of directed graphs to support graph-theoretic interpretations. A cycle cover is a collection of vertex-disjoint directed cycles such that every vertex of $G$ lies on exactly one cycle. A partial cycle cover is similar, but it covers only a subset of the vertices, leaving the rest uncovered. When a cycle cover that includes all vertices is not possible, we are often interested in a maximum partial cycle cover, which is a collection of vertex-disjoint cycles that covers the largest possible number of vertices in the graph.
For a symmetric matrix, it is known that the rank equals the number of nonzero eigenvalues, while the nullity corresponds to the number of zero eigenvalues. In this paper, we show that an analogous relationship holds for the permanental roots of a matrix. Specifically, we prove that for non-negative symmetric matrices, positive semi-definite matrices, and balanced signed graphs, the sum of the permanental rank and permanental nullity equals the order of the matrix.
The remainder of the paper is structured as follows. Section 2 introduces the necessary definitions and preliminary results. Section 3 presents our main theorems, establishing the permanental rank-nullity relationship for certain matrix classes and discusses counterexamples where this relation fails. Finally, Section 5 outlines directions for future research.
2 Preliminaries
In this section, we recall some known results that we will use later to prove our main theorems. We start with a trivial result on the coefficients of the permanental polynomial.
Lemma 2.1 ([19]).
Let $A$ be an $n\times n$ matrix, and let $c_{n-k}$ denote the coefficient of $x^{n-k}$ in the permanental polynomial of $A$. Then $c_{n-k}$ is equal to the sum of all $k\times k$ principal subpermanents of $A$. In particular, $c_{n-1}=\mathrm{trace}(A)$ and $c_{0}=\mathrm{per}(A)$.
Next, we recall a few results on positive semi-definite matrices related to their permanents by [20]. For a matrix $A\in\mathbb{C}^{n\times n}$ and an index set $I,J\subseteq[n]$, we use $A[I,J]$ to denote the submatrix of $A$ formed by rows indexed by $I$ and columns indexed by $J$.
Theorem 2.2 ([20]).
Let $A=(a_{ij})$ be a positive semi-definite Hermitian matrix of order $n$. Then the following properties hold:
  1. 1.
    $\operatorname{per}(A)\geq 0$.
  2. 2.
    $a_{ii}a_{jj}\geq a_{ij}^{2}$ for all $i,j$.
  3. 3.
    $\operatorname{per}(A)\geq\prod_{i=1}^{n}a_{ii}$.
  4. 4.
    All principal submatrices of $A$ are also positive semi-definite, hence $\operatorname{per}(A[I,I])\geq 0$ for all index sets $I\subseteq[n]$.
Signed graphs was introduced by Harary [10] and later formalized by Zaslavsky [30]. A signed graph is a pair $(G,\sigma)$, where $G=(V,E)$ is an undirected graph and $\sigma:E\rightarrow\{-1,+1\}$ assigns a sign to each edge. The corresponding signed adjacency matrix $A^{\sigma}\in\mathbb{\{}0,1,-1\}^{n\times n}$ is defined by
$A^{\sigma}(u,v)=\begin{cases}\sigma(\{u,v\})&\text{if }\{u,v\}\in E(G),\\ 0&\text{otherwise}.\end{cases}$
A signed graph $(G,\sigma)$ is said to be balanced if every cycle in the graph contains an even number of negative edges, or equivalently, if the product of the edge signs around every cycle is positive. We use the next result, which gives a simple condition to determine whether a signed graph is balanced.
Proposition 2.3 ([31]).
A signed graph $(G,\sigma)$ is balanced if and only if its signed adjacency matrix $A^{\sigma}$ is diagonally similar to the adjacency matrix $A$ of the corresponding unsigned graph, i.e. $A^{\sigma}=DAD,$ for some diagonal matrix $D=\mathrm{diag}(\pm 1,\ldots,\pm 1)$.
3 Main Results
In this section, we prove a permanental analog of the rank-nullity theorem for three classes of matrices: non-negative symmetric matrices, positive semi-definite matrices, and balanced signed symmetric matrices. In each case, we show that the sum of the permanental rank and the permanental nullity equals the order of the matrix.
For a $n\times n$ matrix $A$ and index sets $I,J\subseteq[n]$, we use $A[I,J]$ to denote the submatrix of $A$ formed by rows indexed by $I$ and columns indexed by $J$. Let $S\subseteq[n]$ be an index set of size $k$. We write $A[S,S]$ to denote the $k\times k$ principal submatrix of $A$ formed by selecting the rows and columns indexed by $S$.
Now, we begin with a general inequality for permanental rank and permanental nullity that holds for all square matrices.
Theorem 3.1.
For any square matrix $A$ of order $n$,
$\rho_{\operatorname{per}}(A)+\eta_{\operatorname{per}}(A)\geq n.$
Proof.
Let the permanental polynomial of $A$ be given by
$\pi(A,x)=\sum_{i=0}^{n}c_{n-i}x^{n-i}$
then by Lemma 2.1, the coefficient of $x^{n-k}$ in $\pi(A,x)$ is
$c_{n-k}=(-1)^{k}\sum_{|S|=k}\operatorname{per}(A[S,S]).$
Assume that $\rho_{\operatorname{per}}(A)=k$. Then by definition, the permanental rank is the largest order of a principal submatrix of $A$ that has a non-zero permanent. Hence, for all subsets $S$ with $|S|>k$, we have $\operatorname{per}(A[S,S])=0.$ This means that the coefficients $c_{n-i}$ for all $i>k$ are equal to zero. Thus,
$c_{n-i}=(-1)^{i}\sum_{|S|=i}\operatorname{per}(A[S,S])=0\quad\text{for all }i>k.$
Consequently, the permanental polynomial reduces to
$\pi(A,x)$ $=\sum_{i=0}^{k}c_{n-i}x^{n-i}$
$=c_{n}x^{n}+\dots+c_{n-k}x^{n-k}$
$=x^{n-k}(c_{n}x^{k}+\dots+c_{n-k}).$
Since $x^{n-k}$ is a factor, the polynomial must have at least $n-k$ roots equal to zero. So, by definition of permanental nullity, we have $\eta_{\operatorname{per}}(A)\geq n-k.$
[…]
This completes the proof. ∎
Example 3.2.
Let $A=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}.$ Then, $\rho_{\operatorname{per}}(A)=1$, $\eta_{\operatorname{per}}(A)=2$, so $\rho_{\operatorname{per}}(A)+\eta_{\operatorname{per}}(A)=3>n=2.$ This shows that the inequality in Theorem 3.1 can be strict in the general case, and equality may fail without further assumptions.
Before we prove the main results, we introduce alternative characterizations of permanental rank and permanental nullity in terms of the associated graphs. These are valid for all non-negative matrices and will be used in the symmetric and signed cases.
Let $A\in\mathbb{R}_{\geq 0}^{n\times n}$ be a non-negative matrix. Since the entries of $A$ are non-negative, there is no cancellation in the computation of the permanent. Therefore, replacing each non-zero entry with $1$ preserves both the permanental rank and the permanental nullity. Hence, it suffices to consider the 00-$1$ matrix with the same support as $A$, which can be interpreted as the adjacency matrix of a directed graph $G$ with $V(G)=[n]\text{ and edge set }E(G)=\{(i,j)\mid A(i,j)\neq 0\}$ , so each non-zero entry $A(i,j)$ corresponds to a directed edge from vertex $i$ to vertex $j$ in $G$.
Proposition 3.3.
Let $A$ be non-negative matrix, then $\rho_{\mathrm{per}}(A)=k$ if and only if $k$ is the maximum number of arcs in $G$ such that every vertex in the induced subgraph on these $k$ arcs has indegree and outdegree at most $1$.
Proof.
Suppose $A$ has $\rho_{\mathrm{per}}(A)=k$, then there exist index sets $I\subseteq[n]$ and $J\subseteq[n]$, each of cardinality $k$, such that the submatrix $B=A[I,J]$ satisfies
$\operatorname{per}(B)\neq 0.$
This implies that there exists at least one bijection $\pi:I\rightarrow J$ such that
$\prod_{i=1}^{k}A(i,\pi(i))\neq 0,$
that is, $A(i,\pi(i))\neq 0$ for all $i\in I$.
We can then define a subgraph $H\subseteq G$ with edge set
$E(H)=\{(i,\pi(i))\mid i\in I\}.$
Since $\pi$ is a bijection, each vertex in $I$ has exactly one outgoing edge in $H$, and each vertex in $J$ has exactly one incoming edge. Thus, the indegree and outdegree of every vertex in the subgraph $H$ is at most 1. Hence, $\rho_{\mathrm{per}}(A)=k$ implies the existence of such an arc set of size $k$.
Conversely, suppose we have a set $F$ of $k$ arcs from $G$ such that each vertex in the resulting subgraph $H=(V,F)$ has indegree and outdegree at most 1. Since each arc in $F\subseteq E$ is directed, let us define
$I=\{i\in V\mid\exists\,j\in V\text{ such that }(i,j)\in F\}\quad\text{and}\quad J=\{j\in V\mid\exists\,i\in V\text{ such that }(i,j)\in F\},$
that is, $I$ is the set of vertices with outdegree 1 in $H$, and $J$ is the set of vertices with indegree 1. Because there are $k$ arcs in $F$, we have $|F|=|I|=|J|=k$.
Now define a function $\pi:I\to J$ by setting $\pi(i)=j$ whenever $(i,j)\in F$. The condition that each vertex has indegree and outdegree at most 1 ensures that $\pi$ is well-defined and injective. Since $|I|=|J|=k$, it follows that $\pi$ is also surjective, and hence a bijection.
Therefore, the arc set $F$ corresponds to a bijection $\pi:I\to J$ such that
$\prod_{i\in I}A(i,\pi(i))\neq 0,$
since every arc $(i,\pi(i))\in F$ corresponds to a nonzero entry in $A$. So the submatrix $B=A[I,J]$ of order $k$ must have a non-zero permanent, and hence $\rho_{\mathrm{per}}(A)\geq k$. Taking the maximum over all such arc sets gives $\rho_{\mathrm{per}}(A)=k$, as claimed.
Let $A$ be a non-negative matrix of order $n$, and let $G$ be its associated directed graph with adjacency matrix equal to the support of $A$. Construct an augmented graph $G^{\prime}_{A}$ by adding a self-loop of weight $-x$ at each vertex of $G$. Then the matrix $A-xI$ corresponds to the weighted adjacency matrix of $G^{\prime}_{A}$, where original edges retain their weights from $A$, and each vertex has an additional loop of weight $-x$.
Proposition 3.4.
The permanental nullity $\eta_{\mathrm{per}}(A)$ equals to the minimum number of vertices covered by self-loops (of weight $-x$) in any cycle cover of $G^{\prime}_{A}$.
Proof.
The permanental polynomial is defined as
$\pi(A,x)=\operatorname{per}(A-xI).$
Each term in this permanent corresponds to a cycle cover of the graph $G^{\prime}_{A}$, where each vertex is either covered by a directed cycle (of length at least 2) or is covered by a self-loop. Then,
$\operatorname{per}(A-xI)=\sum_{\mathcal{C}}\prod_{e\in\mathcal{C}}w(e),$
where the sum runs over all cycle covers $\mathcal{C}$ of $G^{\prime}_{A}$, and $w(e)$ denotes the weight of edge $e$.
Let $k$ be the minimum number of vertices that are covered by self-loops of weight $-x$ in any cycle cover of $G^{\prime}_{A}$. That is, there exists at least one cycle cover in $G^{\prime}_{A}$ that uses exactly $k$ self-loops of weight $-x$, and covers the remaining $n-k$ vertices using proper directed cycles. The permanent expands as a polynomial in the form:
$\operatorname{per}(A-xI)=x^{k}\cdot p_{1}(x)+x^{k+1}\cdot p_{2}(x)+\cdots+x^{n}\cdot p_{r}(x),$
where each $p_{i}(x)\in\mathbb{R}[x]$ collects contributions from cycle covers using exactly $k+i-1$ self-loops of weight $-x$.
A cycle cover having exactly $k$ self-loops of weight $-x$ contributes a monomial of the form $(-x)^{k}\cdot w$, where $w\neq 0$ is the product of the non-zero weights on the proper cycles. This monomial appears in $x^{k}\cdot p_{1}(x)$, and therefore the constant term $p_{1}(0)\neq 0$. Thus, we can factor out $x^{k}$ and write
$\operatorname{per}(A-xI)=x^{k}\cdot p(x),$
where $p(0)\neq 0$. Hence, the multiplicity of the root zero in $\operatorname{per}(A-xI)$ is exactly $k$, which shows that $\eta_{\mathrm{per}}(A)=k$.
Conversely, if $\eta_{\mathrm{per}}(A)=k$, then $x^{k}$ divides $\operatorname{per}(A-xI)$ but $x^{p}$ $(p>k)$ does not. This implies that there exists at least one cycle cover in $G^{\prime}_{A}$ that uses exactly $k$ self-loops of weight $-x$. Therefore, such a cover certifies that $k$ is the minimum number of self-loops of weight $-x$ required in any cycle cover.
Equivalently, the number of vertices covered by self-loops of weight $-x$ in $G^{\prime}_{A}$ is equal to the number of vertices not included in any maximum partial cycle cover of $G$.
3.1 Non-Negative Symmetric Matrices
Together, Propositions 3.3 and 3.4 show that the permanental rank-nullity identity holds if and only if the following two quantities are equal:
  • $\rho_{\mathrm{per}}(A)$: the maximum number of arcs that can be selected in a subgraph of $G$ where no vertex has indegree or outdegree greater than one.
  • $n-\eta_{\mathrm{per}}(A)$: the number of vertices that can be covered by a maximum partial cycle cover in $G$, that is, a collection of vertex-disjoint directed cycles covering as many vertices as possible.
In the case of non-negative symmetric matrices, this equivalence can be established using the symmetry of the associated graph.
Theorem 3.5.
Let $A$ be a non-negative symmetric matrix of order $n\times n$. Then
$\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n.$
Proof.
Suppose that the number of vertices that can be covered by a maximum partial cycle cover in $G$ is exactly $k$. Each of these $k$ vertices lies on a directed cycle and hence has one incoming and one outgoing arc. Therefore, the union of these cycles forms a subgraph of $G$ consisting of $k$ arcs, where each vertex has indegree and outdegree at most one.
Moreover, it is not possible to add any further arcs without violating the condition that no vertex has an indegree or outdegree greater than one. Thus, this set of $k$ arcs forms a maximum arc set with the required degree constraint.
Conversely, suppose there exists a set of $k$ arcs in the directed graph $G$ which induces a subgraph $H$ in which every vertex has indegree and outdegree at most one. Then each connected component of $H$ is either a directed cycle or a directed path.
Every directed cycle in $H$ has an equal number of vertices and arcs. Also, since $G$ is symmetric, for every arc $(u,v)\in E(G)$, the reverse arc $(v,u)$ also belongs to $E(G)$. In particular, for any directed path in $H$, the symmetry of $G$ allows us to pair adjacent vertices and add reverse arcs to form directed 2-cycles. Thus, the path can be transformed into a union of 2-cycles, each covering two vertices, so that the number of vertices covered equals the number of arcs in the path.
Therefore, the set of $k$ arcs gives a structure that covers $k$ vertices using cycles of length at least 2. Hence, the maximum number of arcs under the given constraints equals the maximum number of vertices that can be covered by a partial cycle cover.
By Proposition 3.3, $\rho_{\mathrm{per}}(A)$ is equal to the number of vertices that can be covered by a maximum partial cycle cover in $G$, and by Proposition 3.4, $\eta_{\mathrm{per}}(A)$ is the number of vertices that remain uncovered in any such cover.
Hence, the sum $\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)$ equals the total number of vertices:
$\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n.$
However, the identity does not hold for all symmetric matrices, as shown in the example below.
Example 3.6.
Consider the symmetric matrix $B=\begin{pmatrix}0&0&1&-1\\ 0&0&1&1\\ 1&1&0&1\\ -1&1&1&0\end{pmatrix},$ which has both positive and negative entries.
We find that $\operatorname{per}(B)=0$ and consider the $3\times 3$ principal submatrix corresponding to indices $\{2,3,4\}$, that is, $B=\begin{pmatrix}0&1&1\\ 1&0&1\\ 1&1&0\end{pmatrix}.$ Then, $\operatorname{per}(B)=2\neq 0.$ Thus, $\rho_{\operatorname{per}}(B)=3.$
For the permanental nullity, we obtain the permanental polynomial, $\pi(B,x)=x^{4}+5x^{2}=x^{2}(x^{2}+5).$ The root $x=0$ has multiplicity 2, hence $\eta_{\operatorname{per}}(B)=2.$
Therefore, $\rho_{\operatorname{per}}(B)+\eta_{\operatorname{per}}(B)=3+2=5\neq 4=n.$
3.2 Positive Semi-Definite Matrices
This class of matrices has a well known structure, and we use some known properties of their permanents to carry out the proof.
Lemma 3.7.
Let $A$ be a positive semi-definite matrix with $\rho_{\operatorname{per}}(A)=k$. Then, there exists an index set $S\subseteq[n]$ of size $k$ such that the $k\times k$ principal submatrix $A[S,S]$ satisfies $\operatorname{per}(A[S,S])\neq 0$.
Proof.
By Proposition 3.1 for any square matrix $A$ of order $n$,
$\rho_{\operatorname{per}}(A)+\eta_{\operatorname{per}}(A)\geq n.$
Let $\rho_{\operatorname{per}}(A)=k$. Then there exists a size $k\times k$ submatrix $A[I,J]$ of $A$ such that $\operatorname{per}(A[I,J])\neq 0,$ which implies that for each $i\in I$ there exists a $j\in J$ such that $A(i,j)\neq 0$.
By property (2) of Theorem 2.2, for such $i$ and $j$, we have $a_{ii}\neq 0$ and $a_{jj}\neq 0$. Hence, the submatrix $A[I\cup J,I\cup J]$ has all diagonal entries nonzero. Then, by property (3) of Theorem 2.2, $\operatorname{per}(A[I\cup J,I\cup J])\neq 0.$
If $I\neq J$, then $|I\cup J|>k$, contradicting $\rho_{\operatorname{per}}(A)=k$. Therefore, we must have $I=J=S$ (say), and hence there exists a principal submatrix $A[S,S]$ of order $k$ such that $\operatorname{per}(A[S,S])\neq 0.$ This completes the proof. ∎
Theorem 3.8.
Let $A$ be a positive semi-definite matrix. Then
$\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n.$
Proof.
By Lemma 3.7, there exists a principal submatrix $A[S,S]$ of order $k=\rho_{\operatorname{per}}(A)$ with $\operatorname{per}(A[S,S])\neq 0$. So the permanental polynomial of $A$ is given by, $\pi(A,x)=x^{n-k}\left(c_{n}x^{k}+\cdots+c_{n-k}\right),$ with $c_{n-k}\neq 0$ by property (4) of Theorem 2.2. Thus, the multiplicity of 00 as a root is exactly $n-k$, implying $\eta_{\operatorname{per}}(A)=n-k$.
Hence, $\eta_{\operatorname{per}}(A)+\rho_{\operatorname{per}}(A)=n.$
This result extends naturally to positive semi-definite Hermitian matrices as the argument relies only on the symmetry of matrices, which is preserved under conjugate transposition.
3.3 $(-1,0,1)$-Balanced Symmetric Matrices
Now, we establish the permanental rank-nullity equality for the class of symmetric matrices associated with balanced signed graphs.
Theorem 3.9.
Let $A^{\sigma}$ be a $n\times n$ real symmetric signed matrix corresponding to a balanced signed graph $G$. Then
$\rho_{\mathrm{per}}(A^{\sigma})+\eta_{\mathrm{per}}(A^{\sigma})=n.$
Proof.
By Proposition 2.3, there exists a diagonal matrix $D=\operatorname{diag}(\pm 1,\ldots,\pm 1)$ such that $A^{\sigma}=DAD,$ where $A$ is the adjacency matrix of the underlying unsigned graph $G$, which is symmetric and has non-negative entries.
Now for any index sets $I,J\subseteq[n]$ with $|I|=|J|$, the corresponding submatrix satisfies
$A^{\sigma}[I,J]=D[I,I]\times A[I,J]\times D[J,J],$
where $D[I,I]$ and $D[J,J]$ are diagonal matrices with entries ±$\pm 1$.
Then the permanent satisfies
$\operatorname{per}(A^{\sigma}[I,J])=\operatorname{per}(D[I,I]\times A[I,J]\times D[J,J])=\left(\prod_{i\in I}D_{ii}\right)\left(\prod_{j\in J}D_{jj}\right)\operatorname{per}(A[I,J]),$
which differs from $\operatorname{per}(A[I,J])$ only by a sign. In particular, $\operatorname{per}(A^{\sigma}[I,J])\neq 0$ if and only if $\operatorname{per}(A[I,J])\neq 0$. Therefore, $\rho_{\mathrm{per}}(A^{\sigma})=\rho_{\mathrm{per}}(A).$ For the permanental nullity, observe that
$\pi(A^{\sigma},x)=\operatorname{per}(A^{\sigma}-xI)=\operatorname{per}(DAD-xI)=\operatorname{per}(D(A-xI)D)=\left(\prod_{i=1}^{n}D_{ii}^{2}\right)\operatorname{per}(A-xI),$
which again differs only by a sign. Thus, $\eta_{\mathrm{per}}(A^{\sigma})=\eta_{\mathrm{per}}(A).$ Since $A$ is a non-negative symmetric matrix, by Theorem 3.5, we have
$\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n.$
It follows that $\rho_{\mathrm{per}}(A^{\sigma})+\eta_{\mathrm{per}}(A^{\sigma})=n,$ as claimed.
This shows that the identity holds when the signed graph is balanced. This result also holds for any real symmetric matrix $A\in\mathbb{R}^{n\times n}$ whose associated signed graph is balanced, as the proof depends only on the balance property and symmetry of $A$, not on the specific values of its nonzero entries. The example below shows that the equality may not hold if the graph is only weakly balanced, meaning its vertex set can be partitioned into multiple clusters such that all positive edges lie within clusters and all negative edges occur between clusters.
Example 3.10.
$A^{\sigma}=\begin{pmatrix}0&0&-1&1\\ 0&0&-1&-1\\ -1&-1&0&-1\\ 1&-1&-1&0\\ \end{pmatrix}$ is a symmetric signed adjacency matrix of a weakly balanced graph. We can partition the vertex set into clusters $V_{1}=\{1,4\},V_{2}=\{3\},\text{ and }V_{3}=\{2\}$ such that all positive edges lie within clusters and all negative edges occur between clusters.
Here, $\rho_{\mathrm{per}}(A^{\sigma})=3$, $\eta_{\mathrm{per}}(A^{\sigma})=2,$ and so $\rho_{\mathrm{per}}(A^{\sigma})+\eta_{\mathrm{per}}(A^{\sigma})=5\neq 4=n.$
4 Future Directions
In this paper, we proved that the permanental rank-nullity identity
$\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n$
holds for non-negative symmetric matrices, positive semi-definite matrices, and $(-1,0,1)$-balanced signed matrices. One natural direction for future work is to find necessary and sufficient conditions under which the identity holds for more general classes of matrices. Can the permanental rank and nullity be used to derive structural results about matrices or graphs, and do they interact with other invariants in interesting ways? Developing such analogs may lead to a broader theory parallel to linear algebra.
From a graph-theoretic point of view, it would be interesting to study how permanental rank and nullity have structural implications on specific graph families, such as planar graphs, regular graphs, or graphs with given properties like girth, degree bounds, or chromatic number. Since Laplacian matrices of graphs are positive semi-definite, our result applies to them, and exploring this further might reveal connections with spanning trees, cuts, or resistance distances.
On the algorithmic side, because computing the permanent is #P-complete, it is important to ask whether $\rho_{\mathrm{per}}(A)$ or $\eta_{\mathrm{per}}(A)$ can be efficiently approximated or estimated, at least for special types of matrices or graphs. These directions aim to build a more complete theory of the permanent that parallels the well-known results involving the determinant and algebraic theory.