### v3

# Permanental Analog of the Rank-Nullity Theorem for Symmetric Matrices

Abstract
The rank-nullity theorem is a core result in the study of matrices. The rank of an $n\times n$ matrix is equal to the size of its largest square submatrix with a nonzero determinant; it can be computed in $O(n^{2.37})$ time. For symmetric matrices, the rank equals to the number of nonzero eigenvalues (that is, the nonzero roots of the characteristic polynomial), which implies that the nullity equals the multiplicity of zero eigenvalue. Similar to the rank, the permanental rank of a matrix is the size of the largest square submatrix with nonzero permanent (whose computation is #P-complete), while the permanental nullity is the multiplicity of zero as a root of the permanental polynomial. In this paper, we give a characterization for symmetric $\{0,\pm 1\}$-matrices to hold permanental rank–nullity. We also prove it holds for symmetric nonnegative matrices, positive semidefinite matrices, and adjacency matrices of balanced signed graphs.
1 Introduction
The determinant of an $n\times n$ matrix $A=(a_{ij})$ is defined as
$\det(A)=\sum_{\sigma\in S_{n}}\operatorname{sgn}(\sigma)\prod_{i=1}^{n}a_{i,\sigma(i)},$
where $S_{n}$ is the set of all permutations of $\{1,2,\ldots,n\}$ and $\operatorname{sgn}$ is the sign of the permutation $\sigma$. The permanent is defined in a similar way, but without the $\operatorname{sgn}$ factor:
$\operatorname{per}(A)=\sum_{\sigma\in S_{n}}\prod_{i=1}^{n}a_{i,\sigma(i)}.$
Although the definition differs only by a sign factor, the computational complexity of the determinant and permanent varies significantly. The determinant can be computed in time $\mathcal{O}(n^{2.37})$ [11], while finding the permanent is known to be #$\#P$-complete [24]. The permanent lacks a well established algebraic or geometric interpretation, and it is neither multiplicative nor invariant under linear combinations of rows or columns. As a result, the permanent has received less attention in the early literature, with nearly all known results at the time compiled in the book [18]. In 1979, Valiant [24] proved that computing the permanent is #$\#P$-complete, thus making the permanent a central object of study in computational complexity theory.
The Pólya Permanent Problem [20] investigates conditions under which the permanent of a given matrix can be transformed into the determinant of a modified matrix. This problem is equivalent to twenty-three other combinatorial and graph-theoretic problems [16], including the enumeration of perfect matchings in bipartite graphs. This motivates the search for algebraic frameworks in which the permanent satisfies identities structurally similar to those of the determinant.
Unlike the determinant, the permanent generally lacks multiplicativity, that is, $\operatorname{per}(AB)\neq\operatorname{per}(A)\operatorname{per}(B)$. Marcus and Minc [15] conjectured that multiplicativity holds only when both matrices are products of permutation and diagonal matrices. Beasley [3] later proved this, showing it is the maximal class for which the permanent is multiplicative. Beyond multiplicativity, several determinant inequalities have been adapted for the permanent. Lieb [13] established a reversed version of the classical Fischer inequality for permanents. Specifically, for a block matrix $A=\begin{pmatrix}B&C\\ C^{*}&D\end{pmatrix}\succeq 0,$ where $A$ is positive semidefinite, the inequality $\operatorname{per}(A)\geq\operatorname{per}(B)\operatorname{per}(D)$ holds. This stands in contrast to the determinant case, where the inequality is reversed, that is, $\det(A)\leq\det(B)\det(D)$ under the same conditions. Carlen, Lieb, and Loss also proved a Hadamard type bound, relating the permanent to the product of the row norms [5]. Heuvers, Cummings, and Bhaskara Rao [10] showed that the permanent satisfies an identity analogous to the Cauchy-Binet formula for the determinant.
Analogous to the characteristic polynomial, the permanental polynomial of an $n\times n$ matrix $A$ is defined as
$\pi(A,x)=\operatorname{per}(A-xI),$
where $I$ is the identity matrix of order $n$. Given a graph $G$ with adjacency matrix $A(G)$, its permanental polynomial is $\pi(G,x)=\mathrm{per}(A(G)-xI)$. The multiset of all roots of $\pi(G,x)$, including multiplicities, is called the per-spectrum of $G$. Turner [23] first introduced permanental polynomials in graph theory, and Merris et al. [17] and Kasum et al. [12] later expanded this work to explore their mathematical properties and uses in chemistry. For more results on permanental polynomials and per-spectrum, see [2, 4, 6, 8, 14, 21, 29].
The permanental nullity, $\eta_{\text{per}}(G)$ defined as the multiplicity of zero as a root in $\pi(G,x)$ was first studied by Wu and Zhang [25]. They established its connection to the matching number via the Gallai–Edmonds structure theorem. They provided exact characterizations for graphs with extremal per-nullities $n-2,n-3,n-4,n-5$, and derived a sharp formula for general graphs involving factor-critical components. Their work also characterized graphs with zero permanental nullity and analyzed their behavior on unicyclic graphs, line graphs, and factor-critical graphs.
The concept of permanental rank, denoted by $\rho_{\text{per}}(A)$ for a matrix $A$, was introduced by Yang Yu [26] as a combinatorial analog of the classical matrix rank. Given a matrix $A\in\mathbb{F}^{n\times m}$, $\rho_{\operatorname{per}}(A)$ is defined as the maximum integer $k$ such that $A$ contains a $k\times k$ submatrix whose permanent is nonzero.
Yu established the following fundamental inequality relating permanental rank and classical rank: $\operatorname{rank}(A)\leq 2\rho_{\operatorname{per}}(A),$ which is shown to be tight. Furthermore, over a field of positive characteristic $p>0$, they proved that if a matrix has a column repeated $p$ times, then its permanent vanishes. This leads to an upper bound of the form: $\rho_{\operatorname{per}}(A)\leq(p-1)\operatorname{rank}(A).$ Motivated by the Alon–Jaeger–Tarsi Conjecture [1], they proposed a stronger conjecture: for any nonsingular $n\times n$ matrix $A$, the permanental rank of the matrix $(A\mid A)\in\mathbb{F}^{n\times 2n}$, formed by horizontally concatenating $A$ with itself, satisfies $\rho_{\operatorname{per}}(A\mid A)=n.$ This has been verified for small dimensions and fields of characteristic $p=3$. Additionally, for any integer $l$ and nonsingular matrices $A_{1},A_{2},\dots,A_{\ell}\in\mathbb{F}^{n\times n}$, they proved the following bound:
$\rho_{\operatorname{per}}(A_{1}A_{2}\cdots A_{\ell})\geq\left(1-\frac{2}{\ell}\right)n,$
though the tightness of this inequality is not yet known. In a probabilistic setting, they further showed that for a uniformly random nonsingular matrix $A$ over a finite field, the probability that $\rho_{\operatorname{per}}(A\mid A)=n$ tends to $1$ as $n\to\infty$, thus providing asymptotic support for the conjecture. Fanaï [7] explored the connection of $\rho_{\operatorname{per}}(A)$ with the existence of partial transversals in additive Latin squares, particularly those arising from group-theoretic structures like Cayley tables.
For a symmetric matrix, it is known that the rank equals the number of nonzero eigenvalues, while the nullity corresponds to the number of zero eigenvalues. In this paper, we demonstrate that a similar relationship exists for the permanental roots of a symmetric matrix. Specifically, we prove a permanental analog of the rank–nullity theorem for several classes of symmetric matrices, including nonnegative symmetric matrices, positive semi-definite matrices, and symmetric matrices corresponding to balanced signed graphs. We also provide a necessary and sufficient condition under which the identity $\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n$ holds for arbitrary symmetric matrices.
The remainder of the paper is structured as follows. Section 2 introduces the necessary definitions, notations, and some preliminary results. Section 3 presents our main theorems, establishing the permanental rank-nullity relationship for a few classes of symmetric matrices, and discusses counterexamples where this relation fails. Finally, Section 4 outlines directions for future work.
2 Preliminaries
Signed graphs were introduced by Harary [9] and later formalized by Zaslavsky [27]. A signed graph $G_{\sigma}$ is an undirected graph where each edge $(u,v)$ has a sign $\sigma(u,v)$ which is either positive (+1) or negative (-1). The corresponding signed adjacency matrix $A_{\sigma}\in\mathbb{\{}0,1,-1\}^{n\times n}$ is defined by
$A_{\sigma}(u,v)=\begin{cases}\sigma(u,v)&\text{if }(u,v)\in E(G),\\ 0&\text{otherwise}.\end{cases}$
In a signed graph, a cycle is called positive if it contains an even number of negative edges, and negative if it contains an odd number of negative edges. A signed graph $G_{\sigma}$ is called balanced if every cycle is positive. The following result provides a simple criterion to determine whether a signed graph is balanced.
Lemma 2.1 ([28]).
A signed graph $G_{\sigma}$ is balanced if and only if its signed adjacency matrix $A_{\sigma}$ is diagonally similar to the adjacency matrix $A$ of the corresponding unsigned graph, that is, $A_{\sigma}=DAD,$ for some diagonal matrix $D=\mathrm{diag}(\pm 1,\ldots,\pm 1)$.
For a $n\times n$ matrix $A$ and index sets $I,J\subseteq[n]$, we use $A[I,J]$ to denote the submatrix of $A$ formed by rows indexed by $I$ and columns indexed by $J$. Let $S\subseteq[n]$ be an index set of size $k$. We write $A[S,S]$ to denote the $k\times k$ principal submatrix of $A$ formed by selecting the rows and columns indexed by $S$. We recall a few results on positive semi-definite matrices related to their permanents by [19].
Lemma 2.2 ([19]).
Let $A\in\mathbb{R}^{n\times n}$ be a positive semi-definite symmetric matrix. Then:
  1. 1.
    $\operatorname{per}(A)\geq 0$.
  2. 2.
    $a_{ii}\,a_{jj}\;\geq\;a_{ij}^{2}$ for all $i,j$.
  3. 3.
    $\operatorname{per}(A)\;\geq\;\displaystyle\prod_{i=1}^{n}a_{ii}$.
  4. 4.
    Every principal submatrix of $A$ is positive semi-definite, hence, $\operatorname{per}\bigl{(}A[I,I]\bigr{)}\geq 0$ for all $I\subseteq[n]$.
We now state some results about the coefficients of the permanental polynomial of a matrix.
Lemma 2.3 ([18]).
Let $A\in\mathbb{R}^{n\times n}$ be a matrix. with permanental polynomial
$\pi(A,x)=\sum_{i=0}^{n}b_{i}x^{n-i}.$
Then $b_{0}=1$, and for $1\leq i\leq n$,
$b_{i}=(-1)^{i}\sum_{|S|=i}\operatorname{per}(A[S,S]),$
where the sum is over all principal $i\times i$ submatrices of $A$.
Interestingly, for the adjacency matrix of a signed graph $G_{\sigma}$, the coefficients $b_{i}$ can be expressed in terms of Sachs subgraphs. Sachs subgraph is a subgraph in which every connected component is either an edge or a cycle. For both unsigned and signed graphs, the coefficients of the permanental polynomial can be expressed using sums over Sachs subgraphs, as shown below.
Lemma 2.4 ([22]).
Let $A_{\sigma}$ be the signed adjacency matrix of a signed graph $G_{\sigma}$ on $n$ vertices, and let
$\pi(A_{\sigma},x)=\sum_{i=0}^{n}s_{i}x^{n-i}$
be its permanental polynomial. Then $s_{0}=1$, and for $1\leq i\leq n$,
$s_{i}=(-1)^{i}\sum_{U_{i}}(-1)^{c^{-}(U_{i})}2^{c(U_{i})},$
where the sum is over all signed Sachs subgraphs $U_{i}$ of $G_{\sigma}$ on $i$ vertices, $c(U_{i})$ is the number of cycles in $U_{i}$, and $c^{-}(U_{i})$ is the number of negative cycles in $U_{i}$.
3 Main Results
In this section, we first establish a general inequality for the permanental rank and permanental nullity that holds for all square matrices. We then provide a necessary and sufficient condition under which the permanental rank–nullity identity holds for $\{0,\pm 1\}$-symmetric matrices. Using this condition, we show that the permanental rank–nullity identity holds for nonnegative symmetric matrices and $\{0,\pm 1\}$-balanced matrices. We also prove that the identity holds for positive semi-definite matrices using a different argument based on their structure.
3.1 General Inequality
Theorem 3.1.
For any square matrix $A\in\mathbb{R}^{n\times n}$,
$\rho_{\operatorname{per}}(A)+\eta_{\operatorname{per}}(A)\geq n.$
Proof.
Let the permanental polynomial of $A$ be
$\pi(A,x)=\sum_{i=0}^{n}b_{i}x^{n-i}.$
By Lemma 2.3, for each $1\leq i\leq n$,
$b_{i}=(-1)^{i}\sum_{|S|=i}\operatorname{per}(A[S\mid S]),$
where the sum is over all principal $i\times i$ submatrices of $A$. Assume that $\rho_{\operatorname{per}}(A)=k$. Then every principal submatrix of order greater than $k$ has zero permanent. So, for all $i>k$, the coefficients
$b_{i}=(-1)^{i}\sum_{|S|=i}\operatorname{per}(A[S,S])=0.$
The permanental polynomial simplifies to
$\pi(A,x)$ $=\sum_{i=0}^{k}b_{i}x^{n-i}$
$=b_{0}x^{n}+b_{1}x^{n-1}+\dots+b_{k}x^{n-k}$
$=x^{n-k}(b_{0}x^{k}+b_{1}x^{k-1}+\dots+b_{k}).$
Since $x^{n-k}$ is a factor, the polynomial must have at least $n-k$ roots equal to zero. So, by definition of permanental nullity, we have $\eta_{\operatorname{per}}(A)\geq n-k.$
[…]
The inequality in Theorem 3.1 can be strict without further assumptions as shown in the example below.
Example 3.2.
Let $A=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}.$ Then, $\rho_{\operatorname{per}}(A)=1$, $\eta_{\operatorname{per}}(A)=2$, so $\rho_{\operatorname{per}}(A)+\eta_{\operatorname{per}}(A)=3>n=2.$
3.2 A General Criterion for Symmetric $\{0,\pm 1\}$-Matrices
Let $A\in\{0,\pm 1\}^{n\times n}$ be a symmetric matrix, and let $G_{\sigma}$ be the signed graph with adjacency matrix $A$. By Lemma 2.4, the permanental polynomial of $A_{\sigma}$ is
$\pi(A_{\sigma},x)=\sum_{i=0}^{n}s_{i}x^{n-i},\quad\text{where }s_{i}=(-1)^{i}\sum_{U_{i}}(-1)^{c^{-}(U_{i})}2^{c(U_{i})}.$
Here the sum is over all Sachs subgraphs $U_{i}$ of $G_{\sigma}$ on $i$ vertices, $c(U_{i})$ is the number of cycles in $U_{i}$, and $c^{-}(U_{i})$ is the number of negative cycles. For each $i=0,1,\dots,n$, define
$E_{i}=(-1)^{i}\sum_{\begin{subarray}{c}U_{i}\\ c^{-}(U_{i})\text{ is even}\end{subarray}}2^{c(U_{i})},\qquad O_{i}=(-1)^{i}\sum_{\begin{subarray}{c}U_{i}\\ c^{-}(U_{i})\text{ is odd}\end{subarray}}2^{c(U_{i})}.$
It implies that, $s_{i}=E_{i}-O_{i}.$
Theorem 3.3.
Let $A\in\{0,\pm 1\}^{n\times n}$ be a symmetric matrix, and let $\rho_{\operatorname{per}}(A)=k$. Then
$\rho_{\operatorname{per}}(A)+\eta_{\operatorname{per}}(A)=n\quad\text{if and only if}\quad E_{k}\neq O_{k}.$
Proof.
By Lemma 2.4, the permanental polynomial of $A$ is
$\pi(A,x)=\sum_{i=0}^{n}s_{i}x^{n-i},\quad\text{where }s_{i}=(-1)^{i}\sum_{U_{i}}(-1)^{c^{-}(U_{i})}2^{c(U_{i})}.$
If $\rho_{\mathrm{per}}(A_{\sigma})=k$, we have $s_{i}=0$ for all $i>k$. Hence
$\pi(A,x)=\sum_{i=0}^{k}s_{i}\,x^{n-i}=x^{n-k}\bigl{(}s_{0}x^{k}+s_{1}x^{k-1}+\cdots+s_{k}\bigr{)}.$
The permanental nullity $\eta_{\mathrm{per}}(A)$ is the multiplicity of zero as a root of $\pi(A,x)$, so we have $\eta_{\mathrm{per}}(A)=n-k$ if and only if $s_{k}\neq 0$. Since $s_{k}=E_{k}-O_{k},$ we conclude that $\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n\text{ if and only if }E_{k}\neq O_{k}.$
3.3 Nonnegative Symmetric Matrices
We begin by giving an alternate graph-theoretic definition of the permanental rank, which applies to all nonnegative matrices.
Let $A$ be a nonnegative $n\times n$ matrix. Since all entries are nonnegative, the computation of the permanent involves no cancellation. Hence, replacing each nonzero entry with 1 does not affect the permanental rank. This allows us to associate with $A$ a directed graph $\vec{G}$ with vertex set $[n]=\{1,2,\dots,n\}$, where an arc $i\to j$ exists if and only if $A(i,j)\neq 0$.
Lemma 3.4.
Let $A$ be a nonnegative $n\times n$ matrix. Then $\rho_{\operatorname{per}}(A)=k$ if and only if $k$ is the maximum number of arcs in the associated digraph $\vec{G}$ such that each vertex in the subgraph induced by these $k$ arcs has in-degree and out-degree at most one.
Proof.
Let $\rho_{\mathrm{per}}(A)=k$, then there exist index sets $I\subseteq[n]$ and $J\subseteq[n]$, each of cardinality $k$, such that the submatrix $B=A[I,J]$ satisfies
$\operatorname{per}(B)\neq 0.$
Since $A$ is nonnegative, there exists at least one bijection $\pi:I\rightarrow J$ such that
$\prod_{i=1}^{k}A(i,\pi(i))\neq 0,$
that is, $A(i,\pi(i))\neq 0$ for all $i\in I$. Define a subgraph $H\subseteq\vec{G}$ with edge set
$E(H)=\{(i,\pi(i))\mid i\in I\}.$
Since $\pi$ is a bijection, each vertex in $I$ has exactly one outgoing edge in $H$, and each vertex in $J$ has exactly one incoming edge. Thus, the indegree and outdegree of every vertex in the subgraph $H$ is at most 1. Hence, $\rho_{\operatorname{per}}(A)=k$ implies there exists a subgraph with exactly $k$ arcs such that each vertex in the subgraph has in-degree and out-degree at most one.
Conversely, suppose we have a set $F$ of $k$ arcs from $\vec{G}$ such that each vertex in the resulting subgraph $H=(V,F)$ has indegree and outdegree at most 1. Since each arc in $F$ is directed, let us define
$I=\{i\in V\mid\exists\,j\in V\text{ such that }(i,j)\in F\}\quad\text{,}$
$J=\{j\in V\mid\exists\,i\in V\text{ such that }(i,j)\in F\},$
that is, $I$ is the set of vertices with outdegree 1 in $H$, and $J$ is the set of vertices with indegree 1. Because there are $k$ arcs in $F$, we have $|F|=|I|=|J|=k$.
Now define a function $\pi:I\to J$ by setting $\pi(i)=j$ whenever $(i,j)\in F$. The condition that each vertex has indegree and outdegree at most 1 ensures that $\pi$ is well-defined and injective. Since $|I|=|J|=k$, it follows that $\pi$ is also surjective, and hence a bijection.
Therefore, the arc set $F$ corresponds to a bijection $\pi:I\to J$ such that
$\prod_{i\in I}A(i,\pi(i))\neq 0,$
since every arc $(i,\pi(i))\in F$ corresponds to a nonzero entry in $A$. So the submatrix $B=A[I,J]$ of order $k$ must have a non-zero permanent, and hence $\rho_{\mathrm{per}}(A)\geq k$. Since any such arc set $F$ gives rise to a submatrix with non-zero permanent, the largest number of arcs that form such a subgraph gives the largest size of a submatrix with non-zero permanent. Therefore, $\rho_{\operatorname{per}}(A)=k$.
Given a nonnegative symmetric matrix $A$, we define its associated undirected graph $G$ with vertex set $[n]=\{1,2,\dots,n\}$ and edge set $E(G)=\{\{i,j\}\mid A(i,j)\neq 0\}.$
Lemma 3.5.
Let $A$ be a nonnegative symmetric matrix. If $\rho_{\operatorname{per}}(A)=k$, then $G$ contains a Sachs subgraph on $k$ vertices.
Proof.
Let $\rho_{\operatorname{per}}(A)=k$. By Lemma 3.4, there exists a set $F$ of $k$ arcs in the digraph $\vec{G}$ such that in the subgraph $H=(V,F)$, each vertex has in-degree and out-degree at most one. Therefore, every connected component of $H$ is either a directed cycle or a directed path.
Now interpret these components in the undirected graph $G$ using the symmetry of $A$. A directed cycle of length at least 3 becomes an undirected cycle on the same vertices. A 2-cycle gives an undirected edge, and a directed path gives a set of disjoint edges by selecting alternate arcs. Ignoring directions, we obtain a subgraph of $G$ on $k$ vertices, where each component is either a cycle or an edge.
Thus, $G$ contains a Sachs subgraph on $k$ vertices. ∎
Theorem 3.6.
Let $A$ be a nonnegative symmetric matrix. Then
$\rho_{\operatorname{per}}(A)+\eta_{\operatorname{per}}(A)=n.$
Proof.
Let $\rho_{\mathrm{per}}(A)=k$, then by Lemma 3.5, there exists a Sachs subgraph on $k$ vertices. Since $A$ is nonnegative, every cycle in any Sachs subgraph $U_{k}$ of $G$ is positive, so $c^{-}(U_{k})=0$ for all $U_{k}$. Therefore,
$b_{k}=(-1)^{k}\sum_{U_{k}}2^{c(U_{k})}\neq 0,$
because at least one such subgraph exists. Since $b_{i}=0$ for all $i>k$, the permanental polynomial satisfies
$\pi(A,x)=\sum_{i=0}^{k}b_{i}x^{n-i}=x^{n-k}(b_{0}x^{k}+\dots+b_{k}),$
with $b_{k}\neq 0$, implying that the multiplicity of zero as a root is exactly $n-k$. Hence,
$\eta_{\mathrm{per}}(A)=n-k\quad\Rightarrow\quad\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n.\qed$
The example below shows that the identity $\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n$ does not always hold in general for all symmetric matrices with both positive and negative entries.
Example 3.7.
Consider the symmetric matrix $B=\begin{pmatrix}0&0&1&-1\\ 0&0&1&1\\ 1&1&0&1\\ -1&1&1&0\end{pmatrix}$. We find that $\operatorname{per}(B)=0$ and consider the $3\times 3$ principal submatrix corresponding to indices $\{2,3,4\}$, that is, $B=\begin{pmatrix}0&1&1\\ 1&0&1\\ 1&1&0\end{pmatrix}.$ Then, $\operatorname{per}(B)=2\neq 0.$ Thus, $\rho_{\operatorname{per}}(B)=3.$ For the permanental nullity, we obtain the permanental polynomial, $\pi(B,x)=x^{4}+5x^{2}=x^{2}(x^{2}+5).$ The root $x=0$ has multiplicity 2, hence $\eta_{\operatorname{per}}(B)=2.$ Therefore, $\rho_{\operatorname{per}}(B)+\eta_{\operatorname{per}}(B)=3+2=5\neq 4=n.$
3.4 $\{0,\pm 1\}$-Balanced Symmetric Matrices
In Section 3.2, we established that the permanental rank–nullity identity holds for any symmetric matrix $A$ if and only if $E_{k}\neq O_{k}$, where $k=\rho_{\mathrm{per}}(A)$. We now show that this condition is always satisfied when $A$ corresponds to a balanced signed graph.
Let $A\in\{0,1\}^{n\times n}$ be the nonnegative symmetric matrix obtained by replacing all $-1$ entries in $A_{\sigma}$ with $1$. Then, the graph $G$ with adjacency matrix $A$ is the underlying unsigned graph of $G_{\sigma}$.
Theorem 3.8.
Let $A_{\sigma}\in\{0,\pm 1\}^{n\times n}$ be a symmetric matrix such that the associated signed graph $G_{\sigma}$ is balanced. Then
$\rho_{\operatorname{per}}(A_{\sigma})+\eta_{\operatorname{per}}(A_{\sigma})=n.$
Proof.
By Lemma 2.1, there exists a diagonal matrix $D=\operatorname{diag}(\pm 1,\ldots,\pm 1)$ such that $A_{\sigma}=DAD,$ where $A$ is the adjacency matrix of the underlying unsigned graph $G$, which is symmetric and has nonnegative entries. Now for any index sets $I,J\subseteq[n]$ with $|I|=|J|$, the corresponding submatrix satisfies
$A_{\sigma}[I,J]=D[I,I]\times A[I,J]\times D[J,J],$
where $D[I,I]$ and $D[J,J]$ are diagonal matrices with entries ±$\pm 1$.
Then, the permanent satisfies
$\operatorname{per}(A_{\sigma}[I,J])$ $=\operatorname{per}(D[I,I]\times A[I,J]\times D[J,J])$
$=\left(\prod_{i\in I}D_{ii}\right)\left(\prod_{j\in J}D_{jj}\right)\operatorname{per}(A[I,J]).$
Hence, $\operatorname{per}(A_{\sigma}[I,J])\neq 0$ if and only if $\operatorname{per}(A[I,J])\neq 0$. Therefore, $\rho_{\mathrm{per}}(A_{\sigma})=\rho_{\mathrm{per}}(A).$
Now, let $\rho_{\operatorname{per}}(A_{\sigma})=k$, then $\rho_{\operatorname{per}}(A)=k$. By Lemma 3.5, the underlying graph $G$ contains a Sachs subgraph on $k$ vertices, which implies $E_{k}\neq 0$.
Since $G_{\sigma}$ is balanced, every cycle contains an even number of negative edges, so $c^{-}(U_{k})$ is even for all $U_{k}$. Hence, $O_{k}=0$.
Therefore, $E_{k}\neq O_{k}$, and by Theorem 3.3, we conclude
$\rho_{\operatorname{per}}(A_{\sigma})+\eta_{\operatorname{per}}(A_{\sigma})=n.$
We say that a number has even parity if it is divisible by 2, and odd parity otherwise. A signed graph $G_{\sigma}$ is said to have uniform cycle parity if every cycle in the graph contains the same parity of negative edges, that is, each cycle has either an even number or an odd number of negative edges.
Corollary 3.9.
Let $A_{\sigma}\in\{0,\pm 1\}^{n\times n}$ be a symmetric matrix whose associated signed graph $G_{\sigma}$ has uniform cycle parity. Then
$\rho_{\operatorname{per}}(A_{\sigma})+\eta_{\operatorname{per}}(A_{\sigma})=n.$
Proof.
Let $\rho_{\operatorname{per}}(A_{\sigma})=k$. By Lemma 3.5, the underlying graph $G$ admits a Sachs subgraph on $k$ vertices. If every cycle in $G_{\sigma}$ has even parity, then $E_{k}\neq 0$ and $O_{k}=0$. While if every cycle in $G_{\sigma}$ has odd parity, then $O_{k}\neq 0$ and $E_{k}=0$. In either case $E_{k}\neq O_{k}$. Therefore, by Theorem 3.3,
$\rho_{\operatorname{per}}(A_{\sigma})+\eta_{\operatorname{per}}(A_{\sigma})=n.$
3.5 Positive Semi-Definite Matrices
This class of matrices has a well-known structure, and we use some known properties of their permanents to carry out the proof.
Lemma 3.10.
Let $A\in\mathbb{R}^{n\times n}$ be positive semi-definite, and suppose $\rho_{\operatorname{per}}(A)=k$. Then there exists an index set $S\subseteq[n]$ with $\lvert S\rvert=k$ such that the $k\times k$ principal submatrix $A[S,S]$ has nonzero permanent, that is, $\operatorname{per}(A[S,S])\neq 0$.
Proof.
By Lemma 3.1 for any square matrix $A$ of order $n$,
$\rho_{\operatorname{per}}(A)+\eta_{\operatorname{per}}(A)\geq n.$
Let $\rho_{\operatorname{per}}(A)=k$. Then there exists a size $k\times k$ submatrix $A[I,J]$ of $A$ such that $\operatorname{per}(A[I,J])\neq 0,$ which implies that for each $i\in I$ there exists a $j\in J$ such that $A(i,j)\neq 0$.
By property (2) of Lemma 2.2, for such $i$ and $j$, we have $a_{ii}\neq 0$ and $a_{jj}\neq 0$. Hence, the submatrix $A[I\cup J,I\cup J]$ has all diagonal entries nonzero. Then, by property (3) of Lemma 2.2, $\operatorname{per}(A[I\cup J,I\cup J])\neq 0.$
If $I\neq J$, then $|I\cup J|>k$, contradicting $\rho_{\operatorname{per}}(A)=k$. Therefore, we must have $I=J=S$ (say), and hence there exists a principal submatrix $A[S,S]$ of order $k$ such that $\operatorname{per}(A[S,S])\neq 0.$ This completes the proof. ∎
Theorem 3.11.
Let $A\in\mathbb{R}^{n\times n}$ be positive semi-definite. Then
$\rho_{\operatorname{per}}(A)\;+\;\eta_{\operatorname{per}}(A)\;=\;n.$
Proof.
By Lemma 3.10, there exists a principal submatrix $A[S,S]$ of order $k=\rho_{\operatorname{per}}(A)$ with $\operatorname{per}(A[S,S])\neq 0$. So the permanental polynomial of $A$ is given by, $\pi(A,x)=x^{n-k}\left(b_{0}x^{k}+\cdots+b_{k}\right),$ with $b_{k}\neq 0$ by property (4) of Lemma 2.2. Thus, the multiplicity of 00 as a root is exactly $n-k$, implying $\eta_{\operatorname{per}}(A)=n-k$.
Hence, $\eta_{\operatorname{per}}(A)+\rho_{\operatorname{per}}(A)=n.$
This result extends naturally to positive semi-definite Hermitian matrices as the argument relies only on the symmetry of matrices, which is preserved under conjugate transposition.
4 Future Directions
In this paper, we proved that the permanental rank–nullity identity
$\rho_{\mathrm{per}}(A)+\eta_{\mathrm{per}}(A)=n$
holds for nonnegative symmetric matrices, positive semi-definite matrices, and $(-1,0,1)$-balanced signed matrices. More generally, we showed that the identity holds for any symmetric matrix $A$ if and only if $E_{k}\neq O_{k}$, where $k=\rho_{\mathrm{per}}(A)$. We can extend the condition to any real matrix by letting each entry $A_{ij}$ be the weight on the directed edge $i\to j$, and then comparing the total weights of positive and negative cycles in directed Sachs subgraphs of size $k$. Although this condition extends to all real matrices, it only becomes more meaningful when we can relate the condition $E_{k}\neq O_{k}$ to matrix properties, such as rank, eigenvalue spectrum, permanental spectrum, or other structural patterns. Can the permanental rank and nullity be used to derive structural results about matrices or graphs, and do they interact with other invariants in interesting ways? Developing such analogs may lead to a broader theory parallel to linear algebra. On the algorithmic side, because computing the permanent is #P-complete, it is important to ask whether $\rho_{\mathrm{per}}(A)$ or $\eta_{\mathrm{per}}(A)$ can be efficiently approximated or estimated, at least for special types of matrices or graphs. These directions aim to build a more complete theory of the permanent that parallels the well-known results involving the determinant and algebraic theory.